# Microscale architected materials for elastic wave guiding: Fabrication and dynamic characterization across length and time scales

### v1

@@ Abstract
We present an experimental protocol for the fabrication and characterization of scalable microarchitected elastic waveguides. Using silicon microfabrication techniques, we develop free-standing 2D truss-based architected waveguides with a maximum diameter of 80 mm, unit cells size of $100$ µm, and minimum beam width of $5$ µm, thus achieving scale separation. To characterize elastic wave propagation, we introduce a custom-built scanning optical pump-probe experiment that enables contactless excitation of elastic wave modes and full spatio-temporal reconstruction of wave propagation across hundreds of unit cells with sub-unit cell resolution. Results on periodic architectures show excellent agreement with finite element simulations and equivalent experimental data at larger length scales. Motivated by scalable computational inverse design, we fabricate a specific example of a spatially graded waveguide and demonstrate its ability to guide elastic waves along an arbitrary pre-designed path.
@@ II.1 Fabrication of free-standing Si-based metamaterials
Samples were fabricated using commercially-processed 100 mm diameter Silion-On-Insulator (SOI) wafers. SOI wafers have three layers bonded to each other called the (1) device, (2) buried oxide (BOX), and (3) handle layers (see the first schematic in Fig. 1). Unit cells were etched on the device layer and the BOX layer removed to generate the free-standing architected film. The sample remains attached to the handle layer at the rim, enabling easy handling. Throughout this study, we use SOI wafers with the two different device layer thicknesses, $10$ and $15$ µm, each with a tolerance of ±$\pm 0.5$ µm. All our wafers have a BOX layer of thickness $2.4\pm 0.12$ µm, and handle layer $400\pm 10$ µm. The steps of the fabrication protocol are summarized as follows:
@@
    A photo-mask — designed and developed in house with a direct laser writer — was used to expose the handle layer, using a commercial (Karl Suss MA6) mask aligner. The exposed PR was developed using standard protocol (KOH solution – AZ400K) in a spray developer, thus exposing the region to be etched on the handle layer. The exposed region was etched using Deep Reactive Ion Etching (DRIE), for $60-90$ minutes.
@@
    As a final step, thin aluminum films of thicknesses $20-50$ nm were deposited on the front and back sides of the sample to act as transducers during optically-induced elastic wave excitation (details in Sec. II.2).
@@ II.2 Pump-probe photoacoustic characterization
Figure 3: Overview of the pump-probe experiment. (a) Top view schematic. The section to the left of the sample holder is the probe section – a home-built heterodyne interferometer with the associated electronic components. To the right of the sample holder is the pump section of the experiment, involving a pulsed laser source with optical elements for filtering and alignment. (b) Side view. The photograph on the left shows the sample assembly (in yellow) with the Mitutoyo objective on the top to focus the probe laser beam to the point of measurement, and a second objective at the bottom to focus the pump beam (scale bar: 100 mm). Sample and pump objective are mounted on a series of linear stages for alignment and scanning control. The schematic on the right shows details of the optical path, scanning and alignment stages. Acronyms: M - mirror; ND - neutral density filter; HWP - half-wave plate; QWP - quarter-wave plate; PBS - polarizing beam splitter; RP - reflecting prism; RR - retro-reflector; L - plano-convex lens; AOM - acousto-optic modulator; A - aperture; WP - Wollaston prism; BPD - balanced photodetector; P - periscope assembly; AFG - arbitrary function generator; LIA - lock in amplifier.
@@ II.2.1 Photoacoustic pump excitation
The pulsed laser is reflected onto the sample through a series of mirrors. The beam passes through a commercially-procured barium borate (BBO) crystal (Fig. 3(a)) which, when aligned, generates the second harmonic of the infrared laser. This serves the dual purpose of safety and visibility during alignment as well as fine control of the pump energy density to limit ablation of fragile samples. The BBO crystal’s use was limited to alignment for this study. The beam also passes through a variable neutral density filter (ND) to control the intensity of the incident pulse. The pump beam then passes through a periscope assembly (P2 in Figs. 3(a,b)) mounted on an $x$-$y$-stage (S3 in Fig. 3(b)) to align the pump beam with the sample assembly, and focused on the sample using a 20$\times$ objective lens (OB1 in Fig. 3(b)). The inset in Fig. 3(b) shows a photograph of the sample assembly with the pump and probe microscope objectives below and above the sample, respectively.
@@ II.2.2 Probe: A custom-built heterodyne interferometer
Commercial scanning laser vibrometers have proven to be powerful in studying waves in metamaterials [62, 63, 11, 12]. However, these systems are ineffective at short length and time scales relevant to microarchitected samples. Some of their primary limitations are the range of accessible frequencies, focusing spot size and hence spatial resolution, the trade-off between scanning resolution and field of view, the signal-to-noise ratio for high-sensitivity measurements, and of course procurement costs. We therefore developed a polarization-sensitive heterodyne interferometer (see the schematics in Fig. 3) to measure spatially resolved particle displacements from sub-unit cell (µm) to structural (mm) length scales, with a displacement resolution on the order of 1 nm, across time scales spanning tens of ns to ms.
@@
$I={}$ $I_{r}+I_{s}+2I_{r}I_{s}\cos{\Bigg{(}\omega_{c}t+\dfrac{2\pi}{\lambda}\delta(t)-\varphi_{0}\Bigg{)}},$ (1)
@@
where $I$ is the intensity of an electromagnetic wave, and $\lambda$ its wavelength. $I_{r}$ and $I_{s}$ are the intensities of the reference and sample beams, respectively, $\delta(t)$ is the time-resolved out-of-plane particle displacement at the measurement point, and $\varphi_{0}$ is the phase difference between the reference and sample beams under ambient conditions (i.e., when $\delta=0$). The third term in Eq. 1 shows that the measured intensity $I$ is modulated by the carrier frequency $\omega_{c}$ under the aforementioned ambient conditions. Particle displacements due to wave motion cause modulation of the phase of this term, hence it is only the argument of this oscillating component (at the carrier frequency $\omega_{c}$) that enables its measurement. The reference and sample intensities, $I_{r}$ and $I_{s}$, respectively, can be controlled through the polarization of the reference and signal laser branches of the interferometer, such that the ambient signal (prior to wave propagation) measured at the oscilloscope is a time-harmonic signal at the carrier frequency, centered at zero. During wave propagation, the particle displacement $\delta(t)$ is directly proportional to the change in phase $\varphi_{m}$ of this signal relative to the ambient reference phase $\varphi_{0}$, such that
$\delta(t)=\dfrac{\lambda}{2\pi}\varphi_{m}.$ (2)
@@
Based on the operating principle described above, an interferometer was built with a continuous wave (CW) He-Ne laser (ThorLabs HNL210LB, 20 mW, 633 nm) as the source. Beam intensity is controlled by a variable neutral density filter (ND). The beam first passes through a half-wave plate (HWP) installed before the polarizing beam splitter (PBS1 in Fig. 3(a)). The HWP is carefully aligned to ensure maximum intensity of the interfered beam, thus increasing the signal-to-noise ratio of the carrier signal (third term in Eq. 1). The beam splitter PBS1 splits the beam into the reference and signal branches of the interferometer, with orthogonal polarization relative to each other. The latter ensures minimum back-reflection into the laser head, thus minimizing power fluctuations. The reference branch first passes through a delay line, using a reflecting prism (RP) and a retroreflector (RR). The delay line was installed to ensure that the path length difference of the interferometer is restricted to below the coherence length of the laser source (on the order of 10 cm for He-Ne lasers). A crucial component of the reference branch is an acousto-optic modulator (AOM), which shifts the optical laser frequency by 80 MHz. The AOM (EQ Photonics GmbH, model# 3080-120) was driven by an RF signal generator (Gooch & Housego, model# 3910), which was triggered by an arbitrary function generator (AFG – Teledyne T3AFG40). An aperture was used to selectively transmit the first-order diffracted output from the AOM, frequency-shifted by 80 MHz (carrier frequency).
The signal branch was collimated using two plano-convex lenses (L1 and L2 in Fig. 3(a)) and directed to the sample using a periscope assembly (P1). This assembly is shown in the side view schematic (Fig. 3(b)). After passing through this mirror assembly, the signal beam is focused on the sample through an objective (OB1 in Fig. 3(b)). The choice of the objective depends on the smallest feature size of the sample—for the experiments in the current study, 20$\times$ or 50$\times$ objectives were used to focus the beam down to spot sizes smaller than 10 µm (the smallest feature size in our samples). The beam was reflected back on the sample path and recombined with the reference branch at the polarizing beam splitter PBS2 (Fig. 3(a)). The back-reflected signal beam requires its polarization to be rotated by 90 for recombination at PBS2, which is achieved by a quarter-wave plate (QWP in Fig. 3(b)).
@@ III.1 Elastic waves in periodic architectures
Alternative representations are shown in Fig. 5. The plot on the left shows the particle displacement of each measurement point (offset along the $y$-axis for visualization), as a function of time. Four representative data sets (indicated by the four colors) are shown with more details on the right. The latter compare the windowed particle displacement data along with the Fourier spectra of the raw data (in gray) and the windowed data (in blue). The excitation signal (data at position 0 µm) shows a rise time on the order of 100 ns. At different positions, Fourier data show qualitative changes (relative to the excitation signal) up to $\sim 4$ MHz (indicated as dashed red lines in the Fourier spectra). This is noticeable at positions $500$ and $1000$ µm. Although the displacement resolution is $<1$ nm, data beyond 10 MHz is not measurable using the current configuration of the interferometer and its electronics (indicated by the gray region in the Fourier spectra). Improved characterization and control of the resolution limits of the interferometer for future high-precision measurements are underway.
@@
To extract wave dispersion information from the data, a 2D Fourier transform was performed on the spatio-temporal data set (represented in Figs. 4(b) and 5). The Fourier spectrum of each time-resolved signal $\hat{u}(f;x=x_{i})$ was normalized against that of the excitation point $\hat{u}_{0}=\hat{u}(f;x=x_{0})$. The resulting data shows only a section of the 2D Brillouin zone due to scanning only along one spatial direction. Multiple line scans were not necessary for this dataset due to the periodicity of the architecture implying symmetry of the Brillouin zone (schematically shown in Fig. 6(a)). Fig. 6(b) presents a color map of the thus-obtained data in the frequency ($f$) vs. (normalized) wave number ($k$) space. The wave number is normalized by the length of a unit cell, $L=100$ µm. The vertical axis on the right indicates the normalized frequency $\bar{f}=\dfrac{fL}{c}$ after [8, 64], where $c\approx 8500$ m/s is the longitudinal wave speed of the material (in this case single-crystal silicon [65]). These normalized metrics allow for a direct comparison of wave dispersion in structures across length and time scales. The solid lines represent dispersion relation data from fully resolved finite element calculations (see Appendix C for simulation details).
Fig. 6(c) from Telgen et al. [64] shows equivalent data for a macroscale metamaterial (dimensions $40\times 40$ cm, unit cell size $5$ mm, beam thickness $500$ µm) machined from aluminum. In their experiments, full 2D data were collected, hence probing $k$-space along two directions. The range of the horizontal axis in Fig. 6(b) is equivalent to the path $\Gamma-X$ in Fig. 6(c). Lines and symbols represent computed dispersion relations from finite element calculations (referring to beam and solid finite elements, respectively).
@@
The match between finite element results and microscale experiments is excellent (see Fig. 6(b)), showing the accuracy of the new microscale wave guides and the pump-probe experiment. However, the data shows an additional intermediate mode between the expected modes. In comparison with finite element data (green lines in Fig. 6(b)), this matches a wave mode that is expected to propagate along the $X-S$-path in $k$-space (see Fig. 6(b)), which is perpendicular to the direction probed in our microscale experiments. Although the exact mechanisms for appearance of this mode are currently unknown, we hypothesize that this could occur either from scattering of waves from defects (e.g., damaged beams or dust particles) in the vicinity of the scan path, or from potential out-of-plane curvature of the microfabricated film. Careful control of the pump pulse and etching parameters is expected to limit these effects.
@@ III.2 Elastic wave guiding in spatially-graded architectures
The inset in Fig. 7(a) shows high-resolution micrographs of individual unit cells in the spatially graded samples (each scale bar 25 µm). The design involves smooth gradients between the shown unit cells, with a uniform beam thickness of $5$ µm and unit cell sizes of $100$ µm. The choice of the beam diameter had two reasons: (1) to test the limits of our fabrication protocol, and (2) to maintain the assumption of slender beams (which admits the accurate prediction of dispersion relations via beam-based finite elements; see a detailed discussion and quantitative comparison in [64]). Fig. 7(b) shows a micrograph of one spatially graded sample. As mentioned in Sec. II.2, the back-window was used to excite the acoustic wave at the center of the design for all measurements. Three line scans, labeled L1, L2, and L3, were performed. Arrival and arrest locations of the signal are marked by circles in Fig. 7(b). At each measurement point, 50 time series datasets were averaged to increase the signal-to-noise ratio. The corresponding particle displacement data is shown in Fig. 7(b), where the positions on the vertical axis of all time series data correspond to the vertical positions in the micrograph on the left. Scanning resolution varied between $200$ µm (for L1 and L3) and $100$ µm (for L2). All measurements were taken for a total line length of 20 mm, hence spanning 200 unit cells. Note that the particle displacements in these experiments are much smaller than those in the experiments on periodic samples reported above. This is partly due to a lower pump pulse energy (used to minimize ablation of the aluminum coating during long-time measurements) and the inherent wave guiding ability of the architecture.
Line scan L1 shows relatively large displacements on the order of $10$ nm at the excitation point. The wave propagates along L1 at a group velocity of approximately 275 m/s (approximated from the average slope in the position-time diagram). We notice a gradual drop in signal (more than an order of magnitude) over $\sim 6$ mm (i.e., 60 unit cells away from the excitation point – marked by the red circle in plot L1 of Fig. 7(b)), followed by loss of signal (below the noise floor) after 70 unit cells from the excitation point. The elastic wave does not propagate beyond this range along L1 because the guided wave is redirected in the direction perpendicular to L1 toward L2.

### v2

@@ Abstract
We present an experimental protocol for the fabrication and characterization of scalable microarchitected elastic waveguides. Using silicon microfabrication techniques, we develop free-standing 2D truss-based architected waveguides with a maximum diameter of 80 mm, unit cells size of $100$ µm, and minimum beam width of 55 µm, thus achieving scale separation. To characterize elastic wave propagation, we introduce a custom-built scanning optical pump-probe experiment that enables contactless excitation of elastic wave modes and full spatio-temporal reconstruction of wave propagation across hundreds of unit cells with sub-unit cell resolution. Results on periodic architectures show excellent agreement with finite element simulations and equivalent experimental data at larger length scales. Motivated by scalable computational inverse design, we fabricate a specific example of a spatially graded waveguide and demonstrate its ability to guide elastic waves along an arbitrary pre-designed path.
@@ II.1 Fabrication of free-standing Si-based metamaterials
Samples were fabricated using commercially-processed 100 mm diameter Silion-On-Insulator (SOI) wafers. SOI wafers have three layers bonded to each other called the (1) device, (2) buried oxide (BOX), and (3) handle layers (see the first schematic in Fig. 1). Unit cells were etched on the device layer and the BOX layer removed to generate the free-standing architected film. The sample remains attached to the handle layer at the rim, enabling easy handling. Throughout this study, we use SOI wafers with the two different device layer thicknesses, $10$ and $15$ µm, each with a tolerance of $\pm 0.5$ µm. All our wafers have a BOX layer of thickness $2.4\pm 0.12$ µm, and handle layer $400\pm 10$ µm. The steps of the fabrication protocol are summarized as follows:
@@
    A photo-mask — designed and developed in house with a direct laser writer — was used to expose the handle layer, using a commercial (Karl Suss MA6) mask aligner. The exposed PR was developed using standard protocol (KOH solution – AZ400K) in a spray developer, thus exposing the region to be etched on the handle layer. The exposed region was etched using Deep Reactive Ion Etching (DRIE), for 609060-90 minutes.
@@
    As a final step, thin aluminum films of thicknesses 205020-50 nm were deposited on the front and back sides of the sample to act as transducers during optically-induced elastic wave excitation (details in Sec. II.2).
@@ II.2 Pump-probe photoacoustic characterization
Figure 3: Overview of the pump-probe experiment. (a) Top view schematic. The section to the left of the sample holder is the probe section – a home-built heterodyne interferometer with the associated electronic components. To the right of the sample holder is the pump section of the experiment, involving a pulsed laser source with optical elements for filtering and alignment. (b) Side view. The photograph on the left shows the sample assembly (in yellow) with the Mitutoyo objective on the top to focus the probe laser beam to the point of measurement, and a second objective at the bottom to focus the pump beam (scale bar: 100 mm). Sample and pump objective are mounted on a series of linear stages for alignment and scanning control. The schematic on the right shows details of the optical path, scanning and alignment stages. Acronyms: M - mirror; ND - neutral density filter; HWP - half-wave plate; QWP - quarter-wave plate; PBS - polarizing beam splitter; RP - reflecting prism; $R$ - retro-reflector; L - plano-convex lens; AOM - acousto-optic modulator; A - aperture; WP - Wollaston prism; BPD - balanced photodetector; P - periscope assembly; AFG - arbitrary function generator; LIA - lock in amplifier.
@@ II.2.1 Photoacoustic pump excitation
[…]
@@ II.2.2 Probe: A custom-built heterodyne interferometer
Commercial scanning laser vibrometers have proven to be powerful in studying waves in metamaterials [62, 63, 11, 12]. However, these systems are ineffective at short length and time scales relevant to microarchitected samples. Some of their primary limitations are the range of accessible frequencies, focusing spot size and hence spatial resolution, the trade-off between scanning resolution and field of view, the signal-to-noise ratio for high-sensitivity measurements, and of course procurement costs. We therefore developed a polarization-sensitive heterodyne interferometer (see the schematics in Fig. 3) to measure spatially resolved particle displacements from sub-unit cell (µm) to structural ($m$) length scales, with a displacement resolution on the order of 1 nm, across time scales spanning tens of ns to ms.
@@
[…]
@@
where II is the intensity of an electromagnetic wave, and $\lambda$ its wavelength. $I_{r}$ and $I_{s}$ are the intensities of the reference and sample beams, respectively, $\delta(t)$ is the time-resolved out-of-plane particle displacement at the measurement point, and $\varphi_{0}$ is the phase difference between the reference and sample beams under ambient conditions (i.e., when $\delta=0$). The third term in Eq. 1 shows that the measured intensity II is modulated by the carrier frequency $\omega_{c}$ under the aforementioned ambient conditions. Particle displacements due to wave motion cause modulation of the phase of this term, hence it is only the argument of this oscillating component (at the carrier frequency $\omega_{c}$) that enables its measurement. The reference and sample intensities, $I_{r}$ and $I_{s}$, respectively, can be controlled through the polarization of the reference and signal laser branches of the interferometer, such that the ambient signal (prior to wave propagation) measured at the oscilloscope is a time-harmonic signal at the carrier frequency, centered at zero. During wave propagation, the particle displacement $\delta(t)$ is directly proportional to the change in phase $\varphi_{m}$ of this signal relative to the ambient reference phase $\varphi_{0}$, such that
[…]
@@
Based on the operating principle described above, an interferometer was built with a continuous wave (CW) He-Ne laser (ThorLabs HNL210LB, 20 mW, 633 nm) as the source. Beam intensity is controlled by a variable neutral density filter (ND). The beam first passes through a half-wave plate (HWP) installed before the polarizing beam splitter (PBS1 in Fig. 3(a)). The HWP is carefully aligned to ensure maximum intensity of the interfered beam, thus increasing the signal-to-noise ratio of the carrier signal (third term in Eq. 1). The beam splitter PBS1 splits the beam into the reference and signal branches of the interferometer, with orthogonal polarization relative to each other. The latter ensures minimum back-reflection into the laser head, thus minimizing power fluctuations. The reference branch first passes through a delay line, using a reflecting prism (RP) and a retroreflector ($R$). The delay line was installed to ensure that the path length difference of the interferometer is restricted to below the coherence length of the laser source (on the order of 10 cm for He-Ne lasers). A crucial component of the reference branch is an acousto-optic modulator (AOM), which shifts the optical laser frequency by 80 MHz. The AOM (EQ Photonics GmbH, model# 3080-120) was driven by an RF signal generator (Gooch & Housego, model# 3910), which was triggered by an arbitrary function generator (AFG – Teledyne T3AFG40). An aperture was used to selectively transmit the first-order diffracted output from the AOM, frequency-shifted by 80 MHz (carrier frequency).
[…]
@@ III.1 Elastic waves in periodic architectures
Alternative representations are shown in Fig. 5. The plot on the left shows the particle displacement of each measurement point (offset along the $y$-axis for visualization), as a function of time. Four representative data sets (indicated by the four colors) are shown with more details on the right. The latter compare the windowed particle displacement data along with the Fourier spectra of the raw data (in gray) and the windowed data (in blue). The excitation signal (data at position 0 µm) shows a rise time on the order of 100 ns. At different positions, Fourier data show qualitative changes (relative to the excitation signal) up to $\sim 4$ MHz (indicated as dashed red lines in the Fourier spectra). This is noticeable at positions $500$ and $1000$ µm. Although the displacement resolution is <1<1 nm, data beyond 10 MHz is not measurable using the current configuration of the interferometer and its electronics (indicated by the gray region in the Fourier spectra). Improved characterization and control of the resolution limits of the interferometer for future high-precision measurements are underway.
@@
To extract wave dispersion information from the data, a 2D Fourier transform was performed on the spatio-temporal data set (represented in Figs. 4(b) and 5). The Fourier spectrum of each time-resolved signal $\hat{u}(f;x=x_{i})$ was normalized against that of the excitation point $\hat{u}_{0}=\hat{u}(f;x=x_{0})$. The resulting data shows only a section of the 2D Brillouin zone due to scanning only along one spatial direction. Multiple line scans were not necessary for this dataset due to the periodicity of the architecture implying symmetry of the Brillouin zone (schematically shown in Fig. 6(a)). Fig. 6(b) presents a color map of the thus-obtained data in the frequency ($f$) vs. (normalized) wave number ($k$) space. The wave number is normalized by the length of a unit cell, $L=100$ µm. The vertical axis on the right indicates the normalized frequency $\dfracfLc\bar{f}=\dfrac${fL}{c} after [8, 64], where $c\approx 8500$ m/s is the longitudinal wave speed of the material (in this case single-crystal silicon [65]). These normalized metrics allow for a direct comparison of wave dispersion in structures across length and time scales. The solid lines represent dispersion relation data from fully resolved finite element calculations (see Appendix C for simulation details).
Fig. 6(c) from Telgen et al. [64] shows equivalent data for a macroscale metamaterial (dimensions $40\times 40$ cm, unit cell size 55 mm, beam thickness $500$ µm) machined from aluminum. In their experiments, full 2D data were collected, hence probing $k$-space along two directions. The range of the horizontal axis in Fig. 6(b) is equivalent to the path $\Gamma-X$ in Fig. 6(c). Lines and symbols represent computed dispersion relations from finite element calculations (referring to beam and solid finite elements, respectively).
@@
The match between finite element results and microscale experiments is excellent (see Fig. 6(b)), showing the accuracy of the new microscale wave guides and the pump-probe experiment. However, the data shows an additional intermediate mode between the expected modes. In comparison with finite element data (green lines in Fig. 6(b)), this matches a wave mode that is expected to propagate along the XSX-S-path in $k$-space (see Fig. 6(b)), which is perpendicular to the direction probed in our microscale experiments. Although the exact mechanisms for appearance of this mode are currently unknown, we hypothesize that this could occur either from scattering of waves from defects (e.g., damaged beams or dust particles) in the vicinity of the scan path, or from potential out-of-plane curvature of the microfabricated film. Careful control of the pump pulse and etching parameters is expected to limit these effects.
@@ III.2 Elastic wave guiding in spatially-graded architectures
The inset in Fig. 7(a) shows high-resolution micrographs of individual unit cells in the spatially graded samples (each scale bar 25 µm). The design involves smooth gradients between the shown unit cells, with a uniform beam thickness of 55 µm and unit cell sizes of $100$ µm. The choice of the beam diameter had two reasons: (1) to test the limits of our fabrication protocol, and (2) to maintain the assumption of slender beams (which admits the accurate prediction of dispersion relations via beam-based finite elements; see a detailed discussion and quantitative comparison in [64]). Fig. 7(b) shows a micrograph of one spatially graded sample. As mentioned in Sec. II.2, the back-window was used to excite the acoustic wave at the center of the design for all measurements. Three line scans, labeled L1, L2, and L3, were performed. Arrival and arrest locations of the signal are marked by circles in Fig. 7(b). At each measurement point, 50 time series datasets were averaged to increase the signal-to-noise ratio. The corresponding particle displacement data is shown in Fig. 7(b), where the positions on the vertical axis of all time series data correspond to the vertical positions in the micrograph on the left. Scanning resolution varied between $200$ µm (for L1 and L3) and $100$ µm (for L2). All measurements were taken for a total line length of 20 mm, hence spanning 200 unit cells. Note that the particle displacements in these experiments are much smaller than those in the experiments on periodic samples reported above. This is partly due to a lower pump pulse energy (used to minimize ablation of the aluminum coating during long-time measurements) and the inherent wave guiding ability of the architecture.
Line scan L1 shows relatively large displacements on the order of $10$ nm at the excitation point. The wave propagates along L1 at a group velocity of approximately 275 m/s (approximated from the average slope in the position-time diagram). We notice a gradual drop in signal (more than an order of magnitude) over $\sim 6$ $m$ (i.e., 60 unit cells away from the excitation point – marked by the red circle in plot L1 of Fig. 7(b)), followed by loss of signal (below the noise floor) after 70 unit cells from the excitation point. The elastic wave does not propagate beyond this range along L1 because the guided wave is redirected in the direction perpendicular to L1 toward L2.